Question 5
The following table contains values of the c.d.f. for the random variable X, which has a geometric
distribution with parameter p = 0.4. Choose the option that gives a correct statement about X.

x 1 2 3 4 5 6 ...
F(x) 0.4 0.64 0.784 0.8704 0.9222 0.9533 ...

A q0.8 = 4
B q0.95 = 5
C m = 3.5
D iqr = 0.384
E None of the above

Question 16
The random variable X has a normal distribution with mean 10 and standard deviation 4, and the
random variable Y has a normal distribution with mean 8 and standard deviation 3; X and Y are
independent. Choose the option that is closest to the probability P(X – Y > 0).
A 0.224
B 0.345
C 0.655
D 0.776
E None of the above

Question 22
The numbers of patients admitted to the Accident and Emergency ward of a hospital in a given
six-hour survey period on Friday evenings were monitored over several weeks. A 95% confidence
interval for the average number admitted during the six-hour period on a Friday evening was
calculated as (32, 76). Choose the option that gives the upper limit of a 95% confidence interval
for the average time in minutes between successive admissions during the six-hour survey period.
A 4.74
B 5.33
C 11.25
D 12.67
E None of the above
Answer #1

Ans:

20)F(4)=0.8708

So, option A is correct.

q0.8 = 4

21)

z=(10-8)/sqrt(4^2+3^2)

z=2/5

z=0.4

P(z>0.4)=0.345

Correct option is B.

22)6 hrs=360 minutes

upper limit(in average time in minutes)=360/76=4.74

Correct option is A.